Required information

Problem 25-2A Analysis and computation of payback period, accounting rate of return, and net present value LO P1, P2, P3

[The following information applies to the questions displayed below.]
  
Most Company has an opportunity to invest in one of two new projects. Project Y requires a $335,000 investment for new machinery with a four-year life and no salvage value. Project Z requires a $335,000 investment for new machinery with a three-year life and no salvage value. The two projects yield the following predicted annual results. The company uses straight-line depreciation, and cash flows occur evenly throughout each year. (PV of $1, FV of $1, PVA of $1, and FVA of $1) (Use appropriate factor(s) from the tables provided.)

Project Y Project Z
Sales $ 390,000 $ 312,000
Expenses
Direct materials 54,600 39,000
Direct labor 78,000 46,800
Overhead including depreciation 140,400 140,400
Selling and administrative expenses 28,000 28,000
Total expenses 301,000 254,200
Pretax income 89,000 57,800
Income taxes (36%) 32,040 20,808
Net income $ 56,960 $ 36,992

Problem 25-2A Part 1

Required:
1. Compute each project’s annual expected net cash flows.
Answer #1
Project Y Project Z
Net income 56960 36992
Depreciation expense 83750 111667
Expected net cash flows 140710 148659
Workings:
Project Y Project Z
Depreciation expense =335000/4 =335000/3